OpenStudy (anonymous):

i forgot how to solve systems of equations? 2a + b = 5 a - b = 1

OpenStudy (anonymous):

These equations are perfectly set up for the elimination method. The first equation has a positive b, and the second equation has a negative b. So, add the two equations down, which will eliminate the b and let you solve for a

OpenStudy (ybarrap):

add them and solve for a.

OpenStudy (anonymous):

so 3a - b = 6 then plug zero in as b to find a? answer choices are (-2,1) (2,1) (2,-1) (-2,-1)

OpenStudy (anonymous):

still dont get it

ganeshie8 (ganeshie8):

you can add them vertically cuz b has opposite coefficient already

ganeshie8 (ganeshie8):

2a + b = 5 a - b = 1 ------------ 3a = 6

ganeshie8 (ganeshie8):

divide 3, you have a !

OpenStudy (anonymous):

but the answer choices are (-2,1) (2,1) (2,-1) (-2,-1)

ganeshie8 (ganeshie8):

3a = 6 divide 3 a = 2

ganeshie8 (ganeshie8):

a = 2 -----------------(1)

ganeshie8 (ganeshie8):

plug in a = 2 in your second equation, and solve b your second equaiton :- a - b = 1 put a = 2 2 - b = 1 solve, b = 1 -----------------------------(2)

ganeshie8 (ganeshie8):

so the solution (a, b) = (2, 1)

OpenStudy (anonymous):

a = 2

ganeshie8 (ganeshie8):

does that make some sense

OpenStudy (anonymous):

wen i plug it in i get b = 3 which isn't one of the points

ganeshie8 (ganeshie8):

how did u get b = 3 ? :)

ganeshie8 (ganeshie8):

please show your steps, so we can see whats happening :)

OpenStudy (anonymous):

a - b = 1 2- b = 1 add 2 on both sides b = 3

ganeshie8 (ganeshie8):

^ you need to add -2 both sides

ganeshie8 (ganeshie8):

a - b = 1 2- b = 1 add -2 on both sides -b = -1

ganeshie8 (ganeshie8):

then multiply -1 both sides, b = 1

OpenStudy (anonymous):

oh I get it now. Thanx alot :)

ganeshie8 (ganeshie8):

good :) you wud have noticed that when u added 2 both sides for 2-b=1, you should get 4-b = 3, you should NOT get b = 3